OpenStudy (anonymous):

Larry Klaus withdrew $310.80 simple interest from an account after 4 years. If he invested $1,110.00, how much was his rate of interest?

OpenStudy (anonymous):

I would like to answer this but I don't think I would be correct I got .89 % every year you might want to get a second opinion.

OpenStudy (anonymous):

not 89% but .89%

OpenStudy (anonymous):

The answer choices are 6%, 7%, 5% and 4% @HelpMePlease?

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (anonymous):

see my point exactly just give me a second.

OpenStudy (anonymous):

brb i think i hot it.

OpenStudy (anonymous):

No problem. @HelpMePlease?

OpenStudy (anonymous):

the answer is 7%

OpenStudy (anonymous):

A/(Pt) = r

OpenStudy (anonymous):

310.80/(1110*4) = .07

OpenStudy (agent0smith):

withdrew $310.80 simple interest from an account after 4 years. If he invested $1,110.00, how much was his rate of interest Do it like the others... 4 years, $1110 initial, r is interest rate $1110*4*r = $310.80 Solve for r, then convert it to a percentage.

OpenStudy (anonymous):

Sorry it took me so long to get that easy answer.

OpenStudy (anonymous):

Yay! Thanks! :) @agent0smith @HelpMePlease?

OpenStudy (anonymous):

I would have never gotten it. Don't apologize :) @HelpMePlease?

OpenStudy (agent0smith):

btw the formula i've been using, but haven't written, is A = Prt P is principal/invested amount, r is rate, t is time. A is amount of interest

OpenStudy (anonymous):

do you agree with my answer of .07?